Calculate q, w, ΔU and ΔH for the reversible adiabatic compression of 0.2 mole
of an ideal gas (diatomic) at 27 C from a volume 1 dm3
to a volume of 0.25
dm3
1
Expert's answer
2020-11-09T14:10:11-0500

Q = 0

Cv = 5R/2 = 20.79 J/(mol*K)

γ = 1.4

T1 = 27+273 = 300 K

T1*V1γ-1 = T2*V2γ-1

T2 = 300*10.4/0.250.4 = 522 K

W = - n*Cv*"\Delta"T = -0.2*20.79*(522-300) = -923 J

ΔU = -W = 923 J

Cp = 7R/2 = 29.10 J/(mol*K)

ΔH = n*Cp*"\Delta"T = 0.2*29.10*(522-300) = 1292 J